# Formulation of entropy through work by Carnot machine and direct derivation of law of entropy non-decrease from Kelvin-Planck principle

### v5

@@ III.1 Gyftopoulos-Beretta entropy
$S_{2}-S_{1}=\frac{W_{0\to 2}^{\rm rev}-W_{0\to 1}^{\rm rev}-W_{1\to 2}^{\rm ad/rev}}{T_{o}},$ (28)
where $W_{1\to 2}^{\rm ad/rev}=W_{1\to 0}^{\rm ad/rev}+W_{0\to 2}^{\rm ad/rev}$. Applying the Kelvin-Planck principle to a cyclic process $A_{0}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{1}\stackrel{{\scriptstyle\rm ad}}{{\longrightarrow}}A_{2}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{0}$, we have
$W_{0\to 1}^{\rm rev}+W_{1\to 2}^{\rm ad/rev}+W_{2\to 0}^{\rm rev}\leq 0\leftrightarrow-W_{0\to 1}^{\rm rev}-W_{1\to 2}^{\rm ad/rev}+W_{0\to 2}^{\rm rev}\geq 0.$ (29)
By using Eq. (29) to Eq. (28), we obtain the law of entropy non-decrease $S_{2}-S_{1}\geq 0$. Meanwhile, it may be considered that the entropy in Eq. (27) is dependent on the choice of the reference heat reservoir through the temperature $T_{o}$. In GB approach, it is shown that Eqs. (28) and (29) are in fact independent of $T_{o}$: The GB entropy is shown to be a state quantity without an explicit calculation of the reversible works, and any heat reservoir at any temperature can be used for the definition of entropy as the heat reservoir merely serves as an auxiliary system (see Sec. 7.4 of Ref. [7] for their proof). In contrast, in the present approach, we evaluated the reversible works explicitly by the use of the Carnot machine and showed that the entropy is independent of the choice of the reference heat reservoir (Eq. (10)).
An advantage of the GB entropy Eq. (27) to the Clausius entropy Eq. (2) is that the GB entropy is determined only from the measurement of $W_{0\to 1}^{\rm rev}$ for a reversible process with a heat reservoir at temperature $T_{o}$ and $W_{0\to 1}^{\rm ad/rev}$ for an adiabatic process, with no need to measure heat. But, these two expressions are related through the expression Eq. (10). See Table 1 for the summary of these three expressions.
@@
III.2 Clausius entropy using isothermal heat
@@
where $W_{2^{\prime}\to 1^{\prime}}^{\rm iso/rev}=U_{2^{\prime}}-U_{1^{\prime}}-T_{o}(S_{2^{\prime}}-S_{1^{\prime}})$ coincides with the minus of the Helmholtz free energy difference of the system before and after the isothermal process. By applying the Kelvin-Planck principle Eq. (5) to Eq. (31), we have $W_{\rm cyc}=T_{o}(S_{1^{\prime}}-S_{2^{\prime}})=T_{o}(S_{1}-S_{2})\leq 0$ and thus we derive the law of entropy non-decrease $S_{1}\leq S_{2}$.
The outline of the above derivation is essentially the same as those in Secs. II.2 and II.3. An important difference is that the calculation of the work here is done by using the Helmholtz free energy difference for a reversible isothermal process, while it is done by using the reversible work Eq. (11) for a reversible temperature-changing process with a Carnot machine in Secs. II.2 and II.3. With this process using a Carnot machine, we were able to apply the Kelvin-Planck principle to show the non-decrease of the entropy Eq. (10), which is essentially the same entropy as the Clausius entropy Eq. (2) in terms of non-isothermal heat.
III.3 Characterization of entropy as thermodynamic cost to create a system’s state from a reference state
@@
In light of the discussion so far, we may characterize entropy as the thermodynamic cost to create a system’s state from a reference state. To this end, let us consider a setup in Fig. 3: From a simple system at state $A_{0}$, a compound system at state $A_{1}$ constituted with $M$ subsystems and a simple system at state $A_{1^{\prime}}$ are created, where the internal energies and volumes of the states $A_{1}$ and $A_{1^{\prime}}$ are set to be the same. The thermodynamic cost, the work taken for creating each state by combining reversible adiabatic processes and reversible temperature-changing processes with a Carnot machine (Fig. 1), is calculated as the minus of the extracted work as
@@
where $U_{1}=\sum_{j=1}^{M}U_{1,j}$ and $S_{1}=\sum_{j=1}^{M}S_{1,j}$, and $U_{1}=U_{1}^{\prime}$. Consider a cyclic process $A_{0}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{1}\stackrel{{\scriptstyle\rm ad}}{{\longrightarrow}}A_{1^{\prime}}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{0}$, where during $A_{1}\stackrel{{\scriptstyle\rm ad}}{{\longrightarrow}}A_{1^{\prime}}$ the allowed adiabatic operation is to remove all the insulated walls separating the subsystems so that the compound system at state $A_{1}$ spontaneously changes to the simple system at state $A_{1^{\prime}}$; the work taken for this process is assumed to vanish as $-W_{1\to 1^{\prime}}^{\rm ad}=0$. The application of Kelvin-Planck principle to this cyclic process yields $W_{0\to 1}^{\rm rev}\leq W_{0\to 1^{\prime}}^{\rm rev}$, which can be expressed in terms of the work cost as
@@
Note that this is equivalent to the law of entropy non-decrease $S_{1}\leq S_{1^{\prime}}$ as confirmed from Eqs. (32) and (33) with $U_{1}=U_{1^{\prime}}$. The expression Eq. (34) implies that the spontaneous process from $A_{1}$ to $A_{1^{\prime}}$ occurs only when $A_{1}$ is more costly than $A_{1^{\prime}}$ to create from $A_{0}$ thermodynamically; the low-entropy (or ordered) structure is hard to create, which is easily seen by comparing Eqs. (32) and (33) with $U_{1}=U_{1^{\prime}}$. This may give a characterization of entropy as the thermodynamic cost to create one state from another, which makes it possible to distinguish between two macroscopic systems of the same internal energy and volume.
@@
[…]
@@
[…]
@@
respectively, where we note $U_{1}=U_{1^{\prime}}$. Because of $(x+y)/2\geq\sqrt{xy}$ for $x,y\geq 0$, we have $-W_{0\to 1}^{\rm rev}\geq-W_{0\to 1^{\prime}}^{\rm rev}$.
IV Conclusion
We derived the law of entropy non-decrease Eq. (1) from the Kelvin-Planck principle Eq. (5) for both simple and compound systems without using the Clausius inequality Eq. (3), where the process that utilizes the Carnot machine operating between the system and the heat reservoir played the essential role in the derivation. Our entropy Eq. (10) is equivalent to the Clausius entropy Eq. (2) and the Gyftopoulos-Beretta entropy Eq. (27), establishing a thermodynamic consistency. We also characterized entropy as the thermodynamic cost to create a system’s state from a reference state. We hope that the present derivation elucidates the principle of thermodynamics. It is also of interest to apply the present formulation of entropy to nonequilibrium entropies that determine adiabatic accessibility between nonequilibrium states [8].
@@ IV Conclusion
In the case using a reversible isothermal process, the maximum work, the right-hand side of Eq. (41), is nothing but the Helmholtz free energy difference (sec. III.2). But this is not the only case, and as we see in Eq. (11), the equality is attained also for the reversible temperature-changing process with a Carnot machine.
Acknowledgements.

### v7

@@ 3.1 Gyftopoulos-Beretta entropy
$S_{2}-S_{1}=\frac{W_{0\to 2}^{\rm rev}-W_{0\to 1}^{\rm rev}-(U_{1}-U_{2})}{T_{o}}.$ (28)
Applying the Kelvin-Planck principle to a cyclic process $A_{0}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{1}\stackrel{{\scriptstyle\rm ad}}{{\longrightarrow}}A_{2}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{0}$, we have
$W_{0\to 1}^{\rm rev}+(U_{1}-U_{2})+W_{2\to 0}^{\rm rev}\leq 0\leftrightarrow-W_{0\to 1}^{\rm rev}-(U_{1}-U_{2})+W_{0\to 2}^{\rm rev}\geq 0.$ (29)
By using Eq. (29) to Eq. (28), we obtain the law of entropy non-decrease $S_{2}-S_{1}\geq 0$. Meanwhile, it may be considered that the entropy in Eq. (27) is dependent on the choice of the reference heat reservoir through the temperature $T_{o}$. In the GB approach, it is shown that Eq. (27) is in fact independent of $T_{o}$: The GB entropy is shown to be a state quantity without an explicit calculation of the reversible work, and any heat reservoir at any temperature can be used for the definition of entropy as the heat reservoir merely serves as an auxiliary system (see Sec. 7.4 of Ref. [6] for their proof). In contrast, in the present approach, we evaluated the reversible work explicitly by the use of the Carnot machine and showed that the entropy is independent of the choice of the reference heat reservoir (Eq. (10)).
An advantage of the GB entropy Eq. (27) to the Clausius entropy Eq. (2) is that the GB entropy is determined only from the measurement of $W_{0\to 1}^{\rm rev}$ for a reversible process with a heat reservoir at temperature $T_{o}$ and $W_{0\to 1}^{\rm ad}=U_{0}-U_{1}$ for an adiabatic process, with no need to measure heat. But, these two expressions are related through the expression Eq. (10). See Table 1 for the summary of these three expressions.
@@
3.2 Clausius entropy using isothermal heat
@@
where $W_{2^{\prime}\to 1^{\prime}}^{\rm iso/rev}=U_{2^{\prime}}-U_{1^{\prime}}-T_{o}(S_{2^{\prime}}-S_{1^{\prime}})$, which stands for the work extracted from the reversible isothermal process, coincides with the minus of the Helmholtz free energy difference of the system before and after the isothermal process. By applying the Kelvin-Planck principle Eq. (5) to Eq. (31), we have $W_{\rm cyc}=T_{o}(S_{1^{\prime}}-S_{2^{\prime}})=T_{o}(S_{1}-S_{2})\leq 0$ and thus we derive the law of entropy non-decrease $S_{1}\leq S_{2}$.
The outline of the above derivation is essentially the same as those in Secs. 2.2 and 2.3. An important difference is that the calculation of the work here is done by using the Helmholtz free energy difference for a reversible isothermal process, while it is done by using the reversible work Eq. (11) for a reversible temperature-changing process with a Carnot machine in Secs. 2.2 and 2.3. With this process using a Carnot machine, we were able to apply the Kelvin-Planck principle to show the non-decrease of the entropy Eq. (10), which is essentially the same entropy as the Clausius entropy Eq. (2) in terms of non-isothermal heat.
3.3 Characterization of entropy as extra thermodynamic cost to create nonuniformity
@@
In light of the discussion so far, we may characterize entropy as an extra thermodynamic cost to create nonuniformity in the system. To this end, let us consider a setup in Fig. 3: From a simple system at state $A_{0}$, a compound system at state $A_{1}$ constituted with $M$ subsystems and a simple system at state $A_{1^{\prime}}$ are created, where the internal energies and volumes of the states $A_{1}$ and $A_{1^{\prime}}$ are set to be the same. The thermodynamic cost, the work taken for creating each state by combining reversible adiabatic processes and reversible temperature-changing processes with a Carnot machine (Fig. 1), is calculated as the minus of the extracted work as
@@
where $U_{1}=\sum_{j=1}^{M}U_{1,j}$ and $S_{1}=\sum_{j=1}^{M}S_{1,j}$, and $U_{1}=U_{1}^{\prime}$. Consider a cyclic process $A_{0}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{1}\stackrel{{\scriptstyle\rm ad}}{{\longrightarrow}}A_{1^{\prime}}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{0}$, where during $A_{1}\stackrel{{\scriptstyle\rm ad}}{{\longrightarrow}}A_{1^{\prime}}$ the allowed adiabatic operation is to remove all the insulated walls separating the subsystems so that the compound system at state $A_{1}$ spontaneously changes to the simple system at state $A_{1^{\prime}}$; the work taken for this process is assumed to vanish as $W_{1\to 1^{\prime}}^{\rm ad}=0$. The application of Kelvin-Planck principle to this cyclic process yields $W_{0\to 1}^{\rm rev}\leq W_{0\to 1^{\prime}}^{\rm rev}$, which can be expressed in terms of the work cost as
@@
Note that this is equivalent to the law of entropy non-decrease $S_{1}\leq S_{1^{\prime}}$ as confirmed from Eqs. (32) and (33) with $U_{1}=U_{1^{\prime}}$. The expression Eq. (34) implies that the spontaneous process from $A_{1}$ to $A_{1^{\prime}}$ occurs only when $A_{1}$ is more costly than $A_{1^{\prime}}$ to create from $A_{0}$ thermodynamically; the nonuniform, low-entropy (ordered) structure is hard to create, which is easily seen by comparing Eqs. (32) and (33) with $U_{1}=U_{1^{\prime}}$. This may give a characterization of entropy as an extra thermodynamic cost that needs to be paid to create nonuniformity in the system, which makes it possible to distinguish between two macroscopic systems of the same internal energy and volume.
@@
$S_{1}-S_{0}=\int_{0}^{1}\frac{\delta Q}{T}=\int_{T_{0}}^{T_{1,1}}\frac{C_{V}}{2T}dT+\int_{T_{0}}^{T_{1,2}}\frac{C_{V}}{2T}dT=C_{V}\ln\frac{\sqrt{T_{1,1}T_{1,2}}}{T_{0}},$ (35)
@@
$S_{1^{\prime}}-S_{0}=\int_{0}^{1^{\prime}}\frac{\delta Q}{T}=\int_{T_{0}}^{T_{1^{\prime}}}\frac{C_{V}}{T}dT=C_{V}\ln\frac{T_{1^{\prime}}}{T_{0}}=C_{V}\ln\frac{T_{1,1}+T_{1,2}}{2T_{0}},$ (36)
@@
respectively, where we note $U_{1}=U_{1^{\prime}}$. Because of $(x+y)/2\geq\sqrt{xy}$ for $x,y\geq 0$, we have $W_{0\to 1}^{\rm rev}\geq-W_{0\to 1^{\prime}}^{\rm rev}$.
4 Conclusion
We derived the law of entropy non-decrease Eq. (1) from the Kelvin-Planck principle Eq. (5) for both simple and compound systems without using the Clausius inequality Eq. (3), where the process that utilizes the Carnot machine operating between the system and the heat reservoir played the essential role in the derivation. Our entropy Eq. (10) is equivalent to the Clausius entropy Eq. (2) and the Gyftopoulos-Beretta entropy Eq. (27), establishing a thermodynamic consistency. We also characterized entropy as the extra thermodynamic cost that needs to be paid to create nonuniformity in the system. We hope that the present derivation elucidates the principle of thermodynamics. It is also of interest to apply the present formulation of entropy to nonequilibrium entropies that determine adiabatic accessibility between nonequilibrium states [8].
{acknowledgment}
@@ 4 Conclusion
In the case using a reversible isothermal process, the maximum work, the right-hand side of Eq. (41), is nothing but the Helmholtz free energy difference (Sec. 3.2). But this is not the only case, and as we see in Eq. (11), the equality is non-trivially attained also for the reversible temperature-changing process with a Carnot machine.